OpenStudy (anonymous):

Is this an exponential growth problem? a population of guppies quadruples every month. How many guppies will there be after 6 months if the beginning population is only 6 guppies? Hi!! I'm taking the exponents and exponential functions test for any of you in connections I just get really psyched out in tests if any of you could help walk me through and confirm the questions that would be so helpful!!

OpenStudy (anonymous):

madriannal have you done an exponential problem before, or is this your first one?

OpenStudy (anonymous):

i have i just want to make sure that this is an exponential growth problem so i can do it if so it would be 6(1+0.03)^6 to solve right?

OpenStudy (anonymous):

nevermind thats not right ugh I'm confused

OpenStudy (anonymous):

P(t) = Po * a^t Let Po be your initial population of 6 guppies. If they quadruple every month, then wouldn't the equation be P(t) = 6 * 4^t at time t = 0, you have 6 at time t = 1 you have 6 * 4 = 24 at time t = 2 you have 6 * 4 * 4 = . . . etc.

OpenStudy (anonymous):

haha that makes so much more sense!

OpenStudy (anonymous):

ok one question is a 3500 principal earns 3% interest compounded semiannually after 20 years what is the balance of the account? I have never heard of semi annually before? not sure how to solve